the following information is provided: c= 0.8 t= 0.15 m= 0.2 calculate the value of the multiplier based on the provided information for the following economies: 1.1.1 closed economy with no government 1.1.2 closed economy with government and tax 1.1 .3 open economy with a government, tax and induced imports
1
Expert's answer
2020-04-23T11:53:30-0400

1.1.1

Y=C+I=0.8+0.2=1

"MPC=\frac{C}{Y}=\frac{0.8}{1}" =0.8

MPS=1-0.8=0.2

"M=\frac{1}{1-MPC}=\frac{1}{0.2}" }=5


1.1.2

let G=0.5

"M=\frac{1}{(1-MPC)(1-t)}\times G=\frac{1}{(1-0.8)(1-0.15)}\times0.5" =2.94


1.1.3

let induced imports=-0.2, f

"M=\frac{1}{(1-c)(1-t)+f}=\frac{1}{(1-0.8)(1-0.15)-0.2}" =2.7